OpenStudy (anonymous):

integrate cosh 2x sinh 3x dx can i solve this using by part?

OpenStudy (solomonzelman):

I would write everything in terms of x.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int \left(\frac{e^{2x}+e^{-2x}}{2}\right)\left(\frac{e^{3x}-e^{-3x}}{2}\right)dx}\)

OpenStudy (anonymous):

why do we need to use definition?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{1}{4}\int \left(e^{2x}+e^{-2x}\right)\left(e^{3x}-e^{-3x}\right)dx}\) \(\color{#000000 }{ \displaystyle \frac{1}{4}\int \left(e^{5x}-e^{-x}+e^x-e^{-5x} \right)dx}\)

OpenStudy (solomonzelman):

that should be simple enough to integrate \(\small \bf ....\)

OpenStudy (solomonzelman):

I think that in this case it is the best, if not the best way.

OpenStudy (anonymous):

i see, what other condition that we need to used definition?

OpenStudy (solomonzelman):

can you rephrase that pleae, because I don't really understand what you want to know...

OpenStudy (anonymous):

in this question, we integrate its definition. so, what type of question that is appropriate to integrate its definition.

OpenStudy (solomonzelman):

we are (rather) integrating the function USING the definition. We know that, \(\color{#000000 }{ \displaystyle \frac{d}{dx} \cosh(x)=\sinh(x) }\)\(\tiny{\\[3.1em]}\) \(\color{#000000 }{ \displaystyle \frac{d}{dx} \sinh(x)=\cosh(x) }\) And these also come from the definition of hyperbolic functions, and by definition I am refering to how they are written in terms of x....

OpenStudy (solomonzelman):

You can do the integration by parts perhaps (with some u-substitution possibly)

OpenStudy (anonymous):

i see, thank you

OpenStudy (solomonzelman):

Although, when it is easier to use the "definition", then do that ... :)

OpenStudy (solomonzelman):

YW